2C3H6O2 + 7 O2 → 6 CO2 + 6 H2O

\bigtriangleupH = -1591 kJ

If you have 2.63 g of the organic fuel (MW = 74.09 g/mol) and 5.63 g of oxygen gas, how many grams of carbon dioxide will you produce?
Answer #1

Molar mass of C3H6O2 = 74.09 g/mol

mass(C3H6O2)= 2.63 g

use:

number of mol of C3H6O2,

n = mass of C3H6O2/molar mass of C3H6O2

=(2.63 g)/(74.09 g/mol)

= 3.55*10^-2 mol

Molar mass of O2 = 32 g/mol

mass(O2)= 5.63 g

use:

number of mol of O2,

n = mass of O2/molar mass of O2

=(5.63 g)/(32 g/mol)

= 0.1759 mol

Balanced chemical equation is:

2 C3H6O2 + 7 O2 ---> 6 CO2 + 6 H2O

2 mol of C3H6O2 reacts with 7 mol of O2

for 3.55*10^-2 mol of C3H6O2, 0.1243 mol of O2 is required

But we have 0.1759 mol of O2

so, C3H6O2 is limiting reagent

we will use C3H6O2 in further calculation

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

According to balanced equation

mol of CO2 formed = (6/2)* moles of C3H6O2

= (6/2)*3.55*10^-2

= 0.1065 mol

use:

mass of CO2 = number of mol * molar mass

= 0.1065*44.01

= 4.687 g

Answer: 4.69 g